# Arboretum: A Large Multimodal Dataset Enabling AI for Biodiversity

### v1

@@ Abstract
We introduce Arboretum, the largest publicly accessible dataset designed to advance AI for biodiversity applications. This dataset, curated from the iNaturalist community science platform and vetted by domain experts to ensure accuracy, includes 134.6 million images, surpassing existing datasets in scale by an order of magnitude. The dataset encompasses image-language paired data for a diverse set of species from birds (Aves), spiders/ticks/mites (Arachnida), insects (Insecta), plants (Plantae), fungus/mushrooms (Fungi), snails (Mollusca), and snakes/lizards (Reptilia), making it a valuable resource for multimodal vision-language AI models for biodiversity assessment and agriculture research. Each image is annotated with scientific names, taxonomic details, and common names, enhancing the robustness of AI model training.
We showcase the value of Arboretum by releasing a suite of CLIP models trained using a subset of 40 million captioned images. We introduce several new benchmarks for rigorous assessment, report accuracy for zero-shot learning, and evaluations across life stages, rare species, confounding species, and various levels of the taxonomic hierarchy.
We anticipate that Arboretum will spur the development of AI models that can enable a variety of digital tools ranging from pest control strategies, crop monitoring, and worldwide biodiversity assessment and environmental conservation. These advancements are critical for ensuring food security, preserving ecosystems, and mitigating the impacts of climate change. Arboretum is publicly available, easily accessible, and ready for immediate use. Please see the project website for links to our data, models, and code.
@@ 1 Introduction
AI advances are poised to play a crucial role in biodiversity conservation, ecology management, and agriculture. Already, AI tools have been shown to enable automated species identification, monitoring of ecological changes, and optimization of crop management [32, 5]. However, standard AI approaches for biodiversity applications persistently face major challenges. Training datasets are labor-intensive and costly to create; they cover only a narrow set of visual concepts; standard vision models excel at single tasks, but require extensive retraining for new tasks; models often struggle with generalizing to unseen labels and new environments, limiting their effectiveness in real-world applications [30, 13]. Models that perform well on benchmarks often fail in the wild [11, 1]. Standard computer vision datasets (ImageNet and its successors) have significant limitations, including incorrectly labeled images, geographical and cultural biases, and overlapping or ill-defined labels, all of which impair the development of high-performant AI models [20]. Consequently, there is a critical need for large, diverse, accurately annotated datasets that are specific to biodiversity, ecology, and agricultural research [23, 19].
@@
Figure 1: The seven taxonomic classes in the Arboretum dataset. Example images of the top 5 most frequent species in each category, including their counts, common names, and scientific names. For clarity, the word ’class’ used in this context is denoted ’taxonomic class’ in the rest of the paper to distinguish it from the other common ML definitions.
In this paper, we make significant contributions to this body of work by curating and releasing Arboretum. This dataset includes 134.6 million captioned images of approximately 326.9K species. The dataset surpasses all existing datasets in scale by an order of magnitude, constituting the largest public, “AI-ready" dataset of curated biodiversity images. The dataset encompasses image-language paired data for a diverse set of species from birds (Aves), spiders/ticks/mites (Arachnida), insects (Insecta), plants (Plantae), fungus/mushrooms (Fungi), snails (Mollusca), and snakes/lizards (Reptilia). See Figure 1 for representative examples, and the project website for details.
Each image in Arboretum is sourced from the iNaturalist community science platform [38] and is annotated with the corresponding common (English) name, scientific (Latin) name, and taxonomic hierarchy. The metadata for each species has been vetted by domain experts, ensuring the accuracy of the text annotations for robust AI model training. We open-source the tooling pipeline used to curate Arboretum, and hope that the community will use this dataset as a valuable resource for further development of multimodal (vision-language) AI models for biodiversity and agriculture research.
To showcase the potential of Arboretum, we further make two technical contributions. First, we train and release ArborCLIP, a suite of vision-language foundation models trained on a subset of approximately 40M Arboretum samples, comprising approximately 33K species. This subset was constructed with standard filtering criteria in order to reduce skewness in species counts. Our models show excellent generalization performance and can support zero-shot (or few-shot) classification with either common or scientific names of unseen taxa. We also anticipate that these models can be fruitfully fine-tuned in the future on datasets for specific biodiversity-related applications.
@@
Second, we rigorously quantify the performance of our foundation models on five existing fine-grained image classification benchmarks, as well on three newly curated test datasets. We find that ArborCLIP models comfortably achieve the state-of-the-art in certain settings, while both the original (OpenAI) CLIP model as well as BioCLIP [35] excel in certain other settings. We analyze these findings in further detail below, but overall we hope that our dataset can be used by the AI community as a testbed for further algorithmic and scaling research in fine-grained image recognition.
@@
Figure 2: Distribution of the Arboretum dataset. (a) Size of each taxonomic class in the Arboretum dataset. (b) Species counts for each taxonomic class. (c) The 40 highest occuring species.
@@
The remainder of this paper is organized as follows. Section 2 introduces the Arboretum dataset, the dataset’s salient characteristics, and a comparison with previous work. Section 3 details our curation methodology. Section 4 introduces our newly proposed test datasets and their characteristics. Section 5 details our new ArborCLIP models and their benchmark performance relative to previous work. Section 6 concludes with a discussion of limitations and potential future directions.
2 The Arboretum Dataset
@@
Figure 3: Treemap diagram of the Arboretum dataset, starting from Kingdom. The nested boxes represent phyla, (taxonomic) classes, orders, and families. Box size represents the relative number of samples.
@@
Arboretum comprises over 134.6M images across seven taxonomic classes —Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. These taxonomic classes were chosen to represent the span of species — outside of charismatic megafauna — which critically impact biodiversity. The images in Arboretum span 326,888 species. Overall, this dataset nearly matches the state-of-the-art curated dataset (TreeOfLife-10M) in terms of species diversity, while comfortably exceeding it in terms of scale by a factor of nearly 13.5×13.5$\times$13.5 ×. Figure 1 shows representative image samples, and Figure 2 displays the distribution of samples according to the seven categories (and the most frequently occurring species). Figure 3 displays phyla, taxonomic classes, orders, and families represented in the dataset.
Each image sample in Arboretum is annotated with a rich amount of curated metadata. This metadata design facilitates easy filtering of species by image count and taxonomic information. The metadata integrates common names, scientific names, and detailed taxonomic hierarchies. For the full list of fields, see Table 1.
Along with the dataset we also release our data curation tooling pipeline, that enables users to easily access and manipulate the dataset. The pipeline allows researchers to select specific categories, visualize data distributions, and manage class imbalance effectively based on their needs. It facilitates the downloading of specific images by their URLs and provides image-text pairs and user-defined chunks to support various AI applications. This pipeline enables users to define different subsets of Arboretum in an easy manner, overall making the dataset AI-ready and easy to use, and reducing the barriers to follow-up research in AI for biodiversity.
@@
We adopt both common and scientific names, since Latin is a low-resource language and current vision-language backbones do not perform well on scientific names alone in a zero-shot manner. We found that a well-structured text description that integrate common names, scientific names, and detailed taxonomic hierarchies facilitates the learning of relationships between Latin and English terms, thereby improving the models’ applicability in scientific contexts [6, 34, 39]. Moreover, incorporating the taxonomic hierarchy enables models to more effectively associate visual data with taxonomic terminology [21, 2]. This matches the guidelines suggested by BioCLIP [35] to enhance model performance and generalization.
Table 1: Annotations provided in the Arboretum Dataset.
@@
Privacy Measures: The images of Arboretum were sourced from the iNaturalist Open Dataset, whose metadata included Personally Identifiable Information (PII). This included information about observers, such as their usernames and sometimes their real names if they have chosen to share that information publicly. We removed all such fields to ensure that no PII is present in the metadata associated with Arboretum samples, ensuring the privacy of all contributors.
Offensive Content: Some of our URLs may point to images that users could find disturbing or harmful, such as photos of dead or dismembered animals. We retained these types of images since they sometimes can provide valuable scientific data about wildlife, including information on predation events, roadkill, and other occurrences relevant to conservation and biodiversity studies. Although iNaturalist relies on user contributions and community moderation to maintain the quality and appropriateness of the data, we acknowledge that the vast and diverse nature of the data means that some offensive or inappropriate content might be present.
Our closest comparisons are with BioScan-1M (which appeared in NeurIPS 2023 Datasets and Benchmarks) and TreeOfLife-10M (which will appear in CVPR 2024). BioScan-1M focuses solely on the Insecta Class and provides scientific names, taxonomic ranks, as well as DNA barcodes. The TreeOfLife-10M dataset comprises 10.4 million images, integrating data from iNat2021 [37], BioScan-1M, and a fresh set of image samples sourced from the Encyclopedia of Life (EOL). It also supports dual-language labels and detailed taxonomic hierarchies, and was used to train the BioCLIP vision-language model. See Table 2 for essential differences.
Table 2: Comparison of Arboretum with existing biodiversity datasets.
Feature Arboretum TreeOfLife BioScan
Size 134.6 million images 10.4 million images 1.1 million images
Diversity 326.9K species 454.1K species 8.3K
@@
Key Features Ready-to-use pipeline, reduce class imbalance, high-quality annotations, supports ArborCLIP Rich hierarchical representations, comprehensive metadata, supports BioCLIP Focus on insects, high-resolution images, detailed taxonomic annotation, DNA codes
@@ 3 Data Collection and Curation Methodology
All of Arboretum is sourced from the iNaturalist Open Data community science platform, which (in all) comprises over 190M biodiversity-relevant observations shared by users. However, there are still significant gaps in usability for AI research. The photos and metadata, although easily downloadable, are provided in four separate metadata sheets that are not ready to use. Taxa information is encoded as numerical IDs, requiring additional API calls and non-trivial lookups to convert these into common or scientific names. The multiple metadata sheets structure is fragmented across four separate files—photos, taxa, observations, and observers—adding complexity to data integration. Managing data balance and filtering out species with too few images can lead to biases towards common (charismatic) species and an imbalanced training process.
Curation of Arboretum.
The iNaturalist Open Dataset comprises a collection of 250.0M images stored on an AWS S3 bucket as of May 27, 2024, with associated metadata in the form of four separate CSV files (indexed as photos, observations, taxa, and observers). Details of each of these files are provided in Section A.5 in the Appendix. While these files contain a plethora of interesitng information, they are designed for rapid lookup, and therefore not “AI-ready". To resolve this, we curate the metadata into a streamlined format for easy usage by AI practitioners.
We first use the corresponding CSV files to populate an SQL database with each CSV file as its own SQL table. We create a new aggregate SQL table by joining photos, observations, and taxa on its relational columns, and discarding columns not relevant to us. We also construct a new column in the aggregated table and populate it with the Amazon S3 URL where the image file is hosted.
We then query this SQL table and extract rows corresponding to the following seven Arboretum categories: Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. [^1] As mentioned above, we choose these 7 classes since they represent non-megafauna that criticially impact biodiversity. Typically, megafauna species are well suited to standard image recognition models, but there remains a pressing need for similar efforts for these particular classes. The iNaturalist Open Data metadata files do not contain common name information; to reconstruct this information, we cross-match species names from the iNaturalist Taxonomy DarwinCore Archive, which is compiled monthly. Once the common name is reconstructed we append this to the SQL table. Finally, we export this table as a CSV file, which we release for public use on HuggingFace.
@@
As described above, Arboretum comprises well-processed metadata with full taxa information and URLs pointing to image files. The metadata can be used to filter specific categories, visualize data distribution, and manage imbalance effectively. We provide a collection of software tools that enable users to easily download, access, and manipulate the dataset. This makes it straightforward for researchers to explore different data subsets.
The citizen science aspect of iNaturalist data leads to a wide variation in the number of observations per species, with some species having only a few records while others have thousands. To mitigate this, our tools can be used to apply user-defined filtering criteria to exclude species with fewer than a specified number of images. Additionally, we imposed an upper limit on the number of images per species to mitigate overfitting.
To address imbalances during training (described further below in our experiments), we download the data and organize into chunked tar files using a semi-global shuffling strategy. Initially, each tar file is shuffled and divided into smaller groups. These groups were then randomly merged to form larger batches, ensuring a balanced distribution of species within each batch. This approach significantly improves dataset integrity by preventing any single species from spanning full batches.

### v2

# BioTrove: A Large Curated Image Dataset Enabling AI for Biodiversity

@@ Abstract
We introduce BioTrove, the largest publicly accessible dataset designed to advance AI applications in biodiversity. Curated from the iNaturalist platform and vetted to include only research-grade data, BioTrove contains 161.9 million images, offering unprecedented scale and diversity from three primary kingdoms: Animalia ("animals"), Fungi ("fungi"), and Plantae ("plants"), spanning approximately 366.6K species. Each image is annotated with scientific names, taxonomic hierarchies, and common names, providing rich metadata to support accurate AI model development across diverse species and ecosystems.
We demonstrate the value of BioTrove by releasing a suite of CLIP models trained using a subset of 40 million captioned images, known as BioTrove-Train. This subset focuses on seven categories within the dataset that are underrepresented in standard image recognition models, selected for their critical role in biodiversity and agriculture: Aves ("birds"), Arachnida ("spiders/ticks/mites"), Insecta ("insects"), Plantae ("plants"), Fungi ("fungi"), Mollusca ("snails"), and Reptilia ("snakes/lizards"). To support rigorous assessment, we introduce several new benchmarks and report model accuracy for zero-shot learning across life stages, rare species, confounding species, and multiple taxonomic levels.
We anticipate that BioTrove will spur the development of AI models capable of supporting digital tools for pest control, crop monitoring, biodiversity assessment, and environmental conservation. These advancements are crucial for ensuring food security, preserving ecosystems, and mitigating the impacts of climate change. BioTrove is publicly available, easily accessible, and ready for immediate use.
@@ 1 Introduction
AI advances are poised to play a crucial role in biodiversity conservation, ecology management, and agriculture. Already, AI tools have been shown to enable automated species identification, monitoring of ecological changes, and optimization of crop management [36, 5]. However, standard AI approaches for biodiversity applications persistently face major challenges. Training datasets are labor-intensive and costly to create; they cover only a narrow set of visual concepts; standard vision models excel at single tasks but require extensive retraining for new tasks; models often struggle with generalizing to unseen labels and new environments, limiting their effectiveness in real-world applications [34, 14]. Models that perform well on benchmarks often fail in the wild [12, 1]. Standard computer vision datasets (ImageNet and its successors) have significant limitations, including incorrectly labeled images, geographical and cultural biases, and overlapping or ill-defined labels, all of which impair the development of high-performant AI models [24]. Consequently, there is a critical need for large, diverse, accurately annotated datasets that are specific to biodiversity, ecology, and agricultural research [27, 23].
@@
Figure 1: Top Seven Phyla in the BioTrove Dataset. This figure displays the seven most frequently occurring phyla within BioTrove, which is curated to include data exclusively from the three primary kingdoms: Animalia, Plantae, and Fungi. For each phylum, the five most common species are shown, including their scientific names, common names, and the number of images per species. The phyla are ordered by species diversity, with the most diverse phylum on the right and the least diverse on the left.
In this paper, we contribute to advancing biodiversity AI research by curating and releasing BioTrove, a dataset comprising 161.9 million captioned images across approximately 366.6K species. This dataset surpasses all previous collections in both scale and diversity, representing the largest publicly available, “AI-ready" dataset of curated biodiversity images. Each image in BioTrove is paired with language data and spans a diverse range of taxonomic groups, including Reptilia (reptiles), Plantae (plants), Mollusca (mollusks), Mammalia (mammals), Insecta (insects), Fungi (fungi), Aves (birds), Arachnida (arachnids), Animalia (animals), Amphibia (amphibians), and Actinopterygii (ray-finned fish). The dataset spans global regions, supporting robust training across diverse environmental contexts. Representative examples are shown in Figure 1, and additional details are provided on the project website.
Each image in BioTrove originates from the iNaturalist community science platform [42] and is annotated with detailed metadata, including the common name, scientific name, and complete taxonomic hierarchy. This curated metadata provides research-grade high-quality text annotations that enhance AI model training. Additionally, we open-source a data management pipeline, BioTrove-Process, to facilitate interaction with BioTrove metadata. With BioTrove-Process, researchers can efficiently filter and balance data by selecting specific taxonomic categories, adjusting for taxonomy level, and managing species distribution to reduce skewness. This enables users to create custom subsets that align with their research goals while maintaining consistency in species representation.
To showcase the capabilities of BioTrove, we introduce two technical contributions. First, we train and release BioTrove-CLIP, a suite of vision-language foundation models, using a subset, BioTrove-Train, consisting of approximately 40M images from BioTrove and representing around 33K species. This subset, constructed with BioTrove-Process, includes diverse taxa, specifically focusing on birds (Aves), spiders/ticks/mites (Arachnida), insects (Insecta), plants (Plantae), fungi (Fungi), snails (Mollusca), and snakes/lizards (Reptilia). These taxonomic classes were selected to capture a broad range of species—outside of charismatic megafauna—that critically impact biodiversity. The models exhibit robust generalization capabilities, demonstrating high zero-shot and few-shot performance on unseen taxa when using either common or scientific names. We anticipate that BioTrove-CLIP will serve as a valuable foundation for biodiversity-related applications and can be further fine-tuned for specific research needs.
@@
Second, we rigorously quantify the performance of our foundation models on five existing fine-grained image classification benchmarks, as well as on three newly curated test datasets. We find that BioTrove-CLIP models comfortably achieve the state-of-the-art in certain settings, while both the original (OpenAI) CLIP model as well as BioCLIP [39] excel in certain other settings. We analyze these findings in further detail below, but overall we hope that our dataset can be used by the AI community as a testbed for further algorithmic and scaling research in fine-grained image recognition.
@@
Figure 2: Distribution of the BioTrove dataset. (a) Size of the top seven Phyla in the BioTrove dataset. (b) Species counts for the top seven Phyla. (c) The 40 highest occurring species in entire BioTrove dataset.
@@
The remainder of this paper is organized as follows. Section 2 introduces the BioTrove dataset, the dataset’s salient characteristics, and a comparison with previous work. Section 3 details our curation methodology. Section 4 introduces our newly proposed test datasets and their characteristics. Section 5 details our new BioTrove-CLIP models and their benchmark performance relative to previous work. Section 6 concludes with a discussion of limitations and potential future directions.
2 The BioTrove Dataset
@@
Figure 3: Treemap diagram of the BioTrove dataset, starting from Kingdom. The nested boxes represent phyla, (taxonomic) classes, orders, and families. Box size represents the relative number of samples.
@@
BioTrove comprises over 161.9 million images spanning 372,966 species. This dataset is an order of magnitude larger than existing biodiversity datasets, such as the state-of-the-art TreeOfLife-10M dataset, which it surpasses in scale by a factor of nearly 13.5×13.5$\times$13.5 × while maintaining comparable species diversity. Figure 1 shows representative image samples, while Figure 2 displays the distribution of samples across the seven major categories with the most frequently observed species. Additionally, Figure 3 illustrates the range of phyla, taxonomic classes, orders, and families represented in the dataset.
BioTrove includes only research-grade data and publicly accessible licensed content for research purposes from iNaturalist, which designates observations as research-grade once they meet strict validation criteria. To qualify, two or more experienced iNaturalist community members—naturalists, biologists, or citizen scientists—must agree on the species identification. Additionally, the observation must meet other requirements, such as a clear photograph and precise geolocation data. Recent experiments have shown that iNaturalist’s Research Grade observations achieve approximately 97% accuracy, underscoring the reliability of this community-based validation process [17]. Furthermore, iNaturalist continuously enhances data quality by refining validator criteria and implementing new data quality assessment measures, ensuring BioTrove remains a robust dataset for scientific use.
Each image sample in BioTrove is enriched with detailed, curated metadata that facilitates efficient filtering by species count and taxonomic information. The metadata includes common names, scientific names, and hierarchical taxonomic data, which enhances the usability of the dataset for AI model training. For the complete list of metadata fields, see Table 1.
Along with the dataset, we also release our data curation tooling pipeline: BioTrove-Process, which enables users to easily access and manipulate the dataset. This pipeline allows researchers to select specific categories across different taxonomic levels, visualize data distributions, and effectively manage class imbalance according to their needs. It facilitates the downloading of specific images by their URLs and provides image-text pairs as well as user-defined chunks to support various AI applications. BioTrove-Process thus enables users to define custom subsets of BioTrove with ease, making the dataset fully AI-ready and reducing barriers to follow-up research in biodiversity-focused AI.
@@
We adopt both common and scientific names since Latin is a low-resource language, and current AI models do not perform well on scientific names alone in a zero-shot manner. We found that a well-structured text description that integrates common names, scientific names, and detailed taxonomic hierarchies facilitates the learning of relationships between Latin and English terms, thereby improving the models’ applicability in scientific contexts [6, 38, 43]. Moreover, incorporating the taxonomic hierarchy enables models to more effectively associate visual data with taxonomic terminology [25, 2]. This matches the guidelines suggested by BioCLIP [39] to enhance model performance and generalization.
Table 1: Annotations provided in the BioTrove Dataset.
@@
Privacy Measures: The images of BioTrove were sourced from the iNaturalist Open Dataset, whose metadata included Personally Identifiable Information (PII). This included information about observers, such as their usernames and sometimes their real names if they have chosen to share that information publicly. We removed all such fields to ensure that no PII is present in the metadata associated with BioTrove samples, ensuring the privacy of all contributors. License: During curation, we took care to include only images from iNaturalist Open Data, which are all licensed under either the CC0, or CC-BY, or CC-BY-NC licenses. This ensures that all our images are available for public research purposes. Offensive Content: Some of our URLs may point to images that users could find disturbing or harmful, such as photos of dead or dismembered animals. We retained these types of images since they sometimes can provide valuable scientific data about wildlife, including information on predation events, roadkill, and other occurrences relevant to conservation and biodiversity studies. Although iNaturalist relies on user contributions and community moderation to maintain the quality and appropriateness of the data, we acknowledge that the vast and diverse nature of the data means that some offensive or inappropriate content might be present.
Our closest comparisons are with BioScan-1M (which appeared in NeurIPS 2023 Datasets and Benchmarks) and TreeOfLife-10M (which will appear in CVPR 2024). BioScan-1M focuses solely on the Insecta Class and provides scientific names, taxonomic ranks, as well as DNA barcodes. The TreeOfLife-10M dataset comprises 10.4 million images, integrating data from iNat2021 [41], BioScan-1M, and a fresh set of image samples sourced from the Encyclopedia of Life (EOL). It also supports dual-language labels and detailed taxonomic hierarchies and was used to train the BioCLIP vision-language model. See Table 2 for essential differences.
Table 2: Comparison of BioTrove with existing biodiversity datasets.
Feature BioTrove TreeOfLife BioScan
Size 161.9 million images 10.4 million images 1.1 million images
Diversity 366.6K species 454.1K species 8.3K
@@
Key Features Ready-to-use pipeline, reduce class imbalance, high-quality annotations, supports BioTrove-CLIP Rich hierarchical representations, comprehensive metadata, supports BioCLIP Focus on insects, high-resolution images, detailed taxonomic annotation, DNA codes
@@ 3 Data Collection and Curation Methodology
All of BioTrove is sourced from the iNaturalist Open Data community science platform, which (in all) comprises over 280M biodiversity-relevant observations shared by users. However, there are still significant gaps in usability for AI research. The photos and metadata, although easily downloadable, are provided in four separate metadata sheets that are not ready to use. Taxa information is encoded as numerical IDs, requiring additional API calls and non-trivial lookups to convert these into common or scientific names. The multiple metadata sheets structure is fragmented across four separate files—photos, taxa, observations, and observers—adding complexity to data integration. Managing data balance and filtering out species with too few images can lead to biases toward common (charismatic) species and an imbalanced training process.
Curation of BioTrove.
The iNaturalist Open Dataset comprises a collection of 284.2 million images stored on an AWS S3 bucket as of 2024-09-27, with associated metadata provided across four separate CSV files (photos, observations, taxa, and observers). Details on each of these files are presented in Section A.5 in the Appendix. Although these files contain a wealth of valuable information, they are structured for rapid retrieval rather than AI-readiness. To address this, we curate the metadata into a streamlined, AI-optimized format.
We populate an SQL database with each CSV file as an individual SQL table, then create an aggregated table by joining photos, observations, and taxa on their relational columns, discarding irrelevant columns. In this aggregate table, we add a new column populated with the Amazon S3 URL and generate individual columns for taxonomic kingdom, phylum, class, order, family, genus, and species.
BioTrove includes only research-grade images from the Animalia, Plantae, and Fungi kingdoms, filtering out other domains to maintain a clear biodiversity focus. To achieve this filtering, we apply strict taxonomic criteria, ensuring only these three kingdoms are represented. The iNaturalist metadata files lack common names, so we reconstruct this information by cross-referencing species names from the iNaturalist Taxonomy DarwinCore Archive, updated monthly. This enriched metadata, including common names, is then appended to the SQL table. The final curated dataset is exported as parquet files, available for public access on HuggingFace.
@@
As outlined, BioTrove includes structured metadata that is both comprehensive and easy to work with, featuring full taxonomic information and direct URLs to image files. To further support accessibility, we release an accompanying software pipeline that allows users to filter specific categories, visualize data distributions, and manage dataset imbalances effectively. These tools make it simple for researchers to interact with BioTrove, creating tailored subsets based on their specific needs. The iNaturalist data, sourced from citizen science contributions, has inherent variability in species representation, with some species documented extensively and others less so. To address this, our tools enable user-defined filters to exclude species with fewer than a set number of images and to cap image counts per species, thus supporting more balanced model training.
To further mitigate dataset imbalances (detailed in our experiments section), we use a semi-global shuffling strategy in which the data is organized into chunked tar files. These files are shuffled, divided into smaller groups, and then merged into larger batches to ensure a balanced species distribution within each batch. This method enhances dataset integrity, helping to prevent the overrepresentation of any single species across the batches.